# Masked Autoencoders are PDE Learners

### v2

@@ Abstract
Neural solvers for partial differential equations (PDEs) have great potential to generate fast and accurate physics solutions, yet their practicality is currently limited by their generalizability. PDEs evolve over broad scales and exhibit diverse behaviors; predicting these phenomena will require learning representations across a wide variety of inputs which may encompass different coefficients, boundary conditions, resolutions, or even equations. As a step towards generalizable PDE modeling, we adapt masked pretraining for physics problems. Through self-supervised learning across PDEs, masked autoencoders can consolidate heterogeneous physics to learn meaningful latent representations and perform latent PDE arithmetic in this space. Furthermore, we demonstrate that masked pretraining can improve PDE coefficient regression and the classification of PDE features. Lastly, conditioning neural solvers on learned latent representations can improve time-stepping and super-resolution performance across a variety of coefficients, discretizations, or boundary conditions, as well as on unseen PDEs. We hope that masked pretraining can emerge as a unifying method across large, unlabeled, and heterogeneous datasets to learn latent physics at scale.
@@ 1 Introduction
The physical world is incredibly complex; physical phenomena can be extremely diverse and span wide spatiotemporal scales—from neuron excitations to turbulent flow to even global climate. Importantly, many of these phenomena can be mathematically modeled with time-dependent partial differential equations (PDEs)[1; 2; 3]. These PDEs are generally analytically intractable and require the use of numerical solvers to obtain approximate solutions. For complex physics, these solutions can often be slow to obtain; furthermore, different PDEs often require a careful design of tailored solvers.
@@
Advances in deep learning in the past decade have led to the design of a new class of solvers for PDEs. These neural solvers can be extremely fast and display resolution invariance; however, neural networks introduce training difficulties and a lack of theoretical guarantees. Many important advances have been made to address these challenges, with SOTA models outperforming numerical solvers on well-studied PDEs under certain setups, as well as proposing error bounds [4; 5; 6; 7; 8; 9; 10].
@@
A current frontier in neural PDE solvers lies in generalizing solvers to different parameters, conditions, or equations, thereby avoiding the need to collect new data and retrain networks when given unseen PDE dynamics. Prior work in this space has explored many methods to achieve this, from directly conditioning on PDE coefficients [11; 12; 13] to pretraining foundation models across various equations [14; 15; 16]. Despite these advances, generalizable neural solvers remain a significant challenge. PDEs can be incredibly diverse, and neural solvers must adapt to different coefficients, geometries, discretizations, or boundary conditions.
@@
As a step towards addressing these challenges, we propose adapting masked pretraining methods to the PDE domain. Specifically, we demonstrate that self-supervised masked pretraining can learn latent structure that can express different coefficients, discretizations, boundary conditions or PDEs under a common representation. Furthermore, we show that masked autoencoders (MAEs) can learn highly structured latent spaces without the need for labels, as well as propose latent physics arithmetic in this space. MAE models can be used to improve downstream tasks such as predicting PDE features or guiding neural solvers in time-stepping or super-resolution through providing meaningful context. Our contributions suggest the possibility to transfer the scalability and flexibility of masked modeling from language and vision domains to physics—creating rich, unified representations of diverse physics through self-supervised learning. We make code available at: https://github.com/anthonyzhou-1/mae-pdes.
@@ 2 Related Work
The field of neural PDE solvers has grown rapidly and has shown great advances in both the accuracy of solutions and the ability to adapt to PDE parameters. Infinite-dimensional neural operators [6; 17; 5] have shown impressive accuracy in solving time-dependent PDEs by learning the mappings between initial conditions and solutions. However, these methods alone have shown brittleness with respect to changing PDE coefficients or boundary conditions [18; 19], prompting recent work to allow neural solvers to adapt to different PDE conditions.
A variety of approaches have considered adding PDE dynamics information to neural solvers. Common neural solvers can support conditional prediction through architecture choices [18], and novel architectures can be designed to explicitly operate with PDE parameter knowledge [8]. Beyond directly conditioning on PDE dynamics, a class of neural PDE solvers has proposed the addition of an encoder or adaptive network to inform a forecaster network of different PDE coefficients [20; 21; 11; 12]. At an even higher level, meta-learning approaches have been adapted to PDE learning to maximize shared learning across different physics [22; 23].
@@
As an effort to work towards more generalizable PDE neural solvers, recent work has followed the success of pretraining and foundational models in the broader deep learning community. Based on contrastive pretraining methods in computer vision problems, [24; 25; 26; 27], contrastive PDE methods aim to leverage equation coefficients [28], physical invariances [29], or Lie point symmetries [30; 31] to define differences in PDE dynamics that can be organized in a latent space. Another approach in PDE pretraining follows observed in-context learning and emergent behavior in LLMs [32; 33; 34] to design neural PDE solvers that are capable of following prompted PDE examples to forecast unseen dynamics [35; 36].
A more straightforward pretraining method focuses on directly training neural solvers to transfer to new PDE dynamics [37; 38; 39]. This approach has also been scaled by training neural solvers with large and diverse training sets to characterize its transfer behavior [14]. As a step toward large-scale modeling, more principled training approaches have been proposed to learn PDE dynamics across diverse physics at scale. Recent work has proposed a combinatorial neural operator that learns different dynamics as separate modules [40], embedding separate PDEs to a common space to do multi-physics prediction [15], incorporating denoising with scalable transformer architectures while training across diverse PDE datasets [16], and using a unified PDE embedding to align LLMs across PDE families [13].
@@
Masked reconstruction is a popular technique popularized by the language processing [41] and vision [42; 43; 44] domains to pretrain models for downstream tasks. Masked modeling is a broad field that spans many masking strategies, architectures, and applications [45]; this ubiquity is attributed to scalability and architecture breakthroughs [46] that allow meaningful context to be learned through masked reconstruction [47]. In the field of neural PDE solvers, masked pretraining has initially been explored as a method to pretrain neural solvers directly [36]. However, we take a separate approach by investigating if models can understand physics through masked self-supervised learning and how these latent representations are manifested. After validating this learning method, we seek to understand if this knowledge can be used to benefit common PDE tasks.
To frame our contribution within these past works, we draw comparisons with broader deep learning efforts to pretrain large encoders through self-supervision, such as BERT [41] or CLIP [48; 49], to be used in downstream tasks. For many reasons (e.g. data availability, architectures, pretraining strategies, compute resources), an equivalent work does not currently exist for PDEs, but we hope to advance this research thrust.
Figure 1: We investigate learning diverse PDE dynamics with masked autoencoders (MAE) and using learned representations to benefit various downstream tasks. (Masked Pretraining) An encoder is trained on unmasked patches of spatiotemporal PDE data, while a decoder reconstructs true data from latent encodings and learned mask tokens. (Supervised Fine-tuning) Pretrained encoders can be used to quickly regress equation coefficients or predict key PDE features. (Conditional Time-stepping) Neural solvers can achieve higher accuracy predictions through conditioning on MAE encodings. (Conditional Super-resolution) SR models can also benefit from conditioning on MAE encodings, using a discretization inversion ($D^{-1}$) and neural operator [50] to predict high-resolution physics.
@@ 3 Methods
Masked Pretraining for PDEs
@@
We adapt the popular Masked Autoencoder (MAE) approach [44; 43; 51] to train ViT models [42; 52] to reconstruct 1D and 2D PDE data. Data are partitioned into non-overlapping patches before a random subset of these patches is sampled to be masked. The masked patches are omitted from the encoder input; the encoder embeds only the unmasked patches through a series of transformer blocks, which allows for larger encoders and faster training at large masking ratios [44]. The encoded patches are then recombined with mask tokens according to their position in the PDE solution. Positional embeddings are added again to preserve positional information before the latent encodings and mask tokens are decoded. The decoder can be shallower and narrower than the encoder because it is discarded in downstream tasks [44], which further reduces training costs. The decoded tokens are projected into the PDE space through a linear layer before reshaping the patches to the input dimension. Lastly, the output is compared to ground truth PDE data through a MSE loss. For additional details on MAE implementation and ablations, we direct readers to Appendix A.1.
Lie Point Symmetry Data Augmentation
@@
To emulate a larger pretraining dataset and aid in generalization, we investigate using Lie augmentations during pretraining, an approach that follows the use of data augmentations in vision or video pretraining [44; 43; 51]. Given a PDE, one can derive its Lie symmetries as a set of transformations $\{g_{1},\dots,g_{i}\}$, each with a variable $\epsilon_{i}$ that modulates the magnitude of the transformation. At training time, we apply $g_{i}$ sequentially, each with a randomly sampled $\epsilon_{i}$ to randomly augment PDE samples. This augmented PDE sample could represent a solution that has been shifted in space, time, or magnitude, among other transformations, but still propagates dynamics according to the original PDE. For a more detailed discussion of Lie point symmetries for PDEs, we refer the reader to Olver (1986) and Mialon et al. (2023). We find that the use of Lie augmentations improves MSE loss when reconstructing a validation set of PDE data; for a detailed comparison, we invite readers to see Appendix A.2.
@@
PDE data are often discretized on a mesh, which can be both irregular and of different resolutions. This can present a challenge for vanilla transformers, which use the same positional encodings regardless of variations in input sequence length; indeed, many approaches have been proposed to adapt ViT models to inputs of varying resolutions or patch sizes [54; 55; 56; 57]. To avoid changes to the ViT backbone, we consider padding variable input sequences to a common length. We also consider an alternate strategy which interpolates positional embeddings to the desired length. Lastly, we investigate encoding the spatial coordinates of the PDE solution grid through an embedder network and incorporating this information by adding an additional token to the input. These modifications tend to improve reconstruction performance when pretraining on multi-resolution PDE data. For additional details and ablation studies, see Appendix A.3.
4 Experiments
4.1 Equations Considered
We describe a variety of PDEs used for masked pretraining and downstream evaluation. In 1D, we pretrain MAE models on the KdV-Burgers equation only, while in 2D we pretrain on the Heat, Advection, and Burgers equations simultaneously. In all PDEs, coefficients and forcing terms are randomly sampled to produce diverse dynamics within a dataset. For additional details regarding PDE parameter sampling, initial conditions, and numerical methods, and we refer readers to Appendix B.

### v3

@@ Abstract
Neural solvers for partial differential equations (PDEs) have great potential to generate fast and accurate physics solutions, yet their practicality is currently limited by their generalizability. PDEs evolve over broad scales and exhibit diverse behaviors; predicting these phenomena will require learning representations across a wide variety of inputs which may encompass different coefficients, boundary conditions, resolutions, or even equations. As a step towards generalizable PDE modeling, we adapt masked pretraining for physics problems. Through self-supervised learning across PDEs, masked autoencoders can consolidate heterogeneous physics to learn rich latent representations. We show that learned representations can generalize to a limited set of unseen equations or parameters and are meaningful enough to regress PDE coefficients or the classify PDE features. Furthermore, conditioning neural solvers on learned latent representations can improve time-stepping and super-resolution performance across a variety of coefficients, discretizations, or boundary conditions, as well as on certain unseen PDEs. We hope that masked pretraining can emerge as a unifying method across large, unlabeled, and heterogeneous datasets to learn latent physics at scale.
@@ 1 Introduction
The physical world is incredibly complex; physical phenomena can be extremely diverse and span wide spatiotemporal scales—from neuron excitations to turbulent flow to even global climate. Importantly, many of these phenomena can be mathematically modeled with time-dependent partial differential equations (PDEs)(FitzHugh, 1961; Nagumo et al., 1962; Lorenz, 1963). These PDEs are generally analytically intractable and require the use of numerical solvers to obtain approximate solutions. For complex physics, these solutions can often be slow to obtain; furthermore, different PDEs often require a careful design of tailored solvers.
@@
Advances in deep learning have led to the design of a new class of solvers for PDEs. These neural solvers can be extremely fast and display resolution invariance; however, neural networks introduce training difficulties and a lack of theoretical guarantees. Many important advances have been made to address these challenges, with models becoming faster than numerical solvers within well-studied PDEs under certain setups, proposing error bounds, and being extended to solve real-world problems. (Raissi et al., 2019; Lu et al., 2019; Li et al., 2020; Cao, 2021; Brandstetter et al., 2022a; Li et al., 2023; Kovachki et al., 2021; Li et al., 2024a).
@@
A current frontier in neural PDE solvers lies in generalizing solvers to different parameters, conditions, or equations, thereby avoiding the need to collect new data and retrain networks when given unseen PDE dynamics. Prior work in this space has explored many methods to achieve this, from directly conditioning on PDE coefficients (Takamoto et al., 2023; Lorsung et al., 2024; Shen et al., 2024) to pretraining foundation models across various equations (Subramanian et al., 2023; McCabe et al., 2023; Hao et al., 2024). Despite these advances, generalizable neural solvers remain a significant challenge. PDEs can be incredibly diverse, and neural solvers must adapt to different coefficients, geometries, discretizations, or boundary conditions.
As a step towards addressing generalizability, we propose adapting masked pretraining methods to the PDE domain. This is motivated by the observation that masked pretraining can learn highly meaningful and broad knowledge in the computer vision and language domains (Devlin et al., 2018). In addition, masked modeling approaches are known to scale well to large and diverse datasets (He et al., 2021). In practice, masked pretraining can also be easily implemented since it makes no prior assumptions about the PDE data, and modern training pipelines and masked modeling architectures are efficient (Vaswani et al., 2023). Lastly, it is hypothesized that masked modeling approaches can generalize well due to having very low inductive bias from masking large portions of the data (Feichtenhofer et al., ).
@@
As such, to study masked modeling within the PDE domain, we train masked autoencoders on a diverse set of 1D and 2D PDE data and evaluate their learned representations. We demonstrate that self-supervised masked pretraining can learn latent structure that can express different coefficients, discretizations, boundary conditions or PDEs under a common representation. Furthermore, we show that masked autoencoders (MAEs) can learn highly structured latent spaces through masking alone. MAE models can be used to improve downstream tasks such as predicting PDE features or guiding neural solvers in time-stepping or super-resolution through providing meaningful context. Our contributions suggest the possibility to transfer the scalability and flexibility of masked modeling from language and vision domains to physics—creating rich, unified representations of diverse physics through self-supervised learning. We provide the code and datasets used in this study here: https://github.com/anthonyzhou-1/mae-pdes .
@@ 2 Related Work
The field of neural PDE solvers has grown rapidly and has shown great advances in both the accuracy of solutions and the ability to adapt to PDE parameters. Infinite-dimensional neural operators (Li et al., 2020; Kovachki et al., 2023; Lu et al., 2019) have shown impressive accuracy in solving time-dependent PDEs by learning the mappings between initial conditions and solutions. However, these methods alone have shown brittleness with respect to changing PDE coefficients or boundary conditions (Gupta and Brandstetter, 2022; Lu et al., 2021), prompting recent work to allow neural solvers to adapt to different PDE conditions.
A variety of approaches have considered adding PDE dynamics information or time-dependent trends to neural solvers. Common neural solvers can support conditional prediction through architecture choices (Gupta and Brandstetter, 2022), and novel architectures can be designed to explicitly operate with PDE parameter knowledge (Brandstetter et al., 2022a). Beyond directly conditioning on PDE dynamics, a class of neural PDE solvers has proposed the addition of an encoder or adaptive network to inform a forecaster network of different PDE coefficients (Wang et al., 2021a; Kirchmeyer et al., ; Takamoto et al., 2023; Lorsung et al., 2024). At an even higher level, meta-learning approaches have been adapted to PDE learning to maximize shared learning across different physics (Yin et al., 2021; Zhang et al., 2023a).
@@
As an effort to work towards more generalizable PDE neural solvers, recent work has followed the success of pretraining and foundational models in the broader deep learning community. Based on contrastive pretraining methods in computer vision problems, (Chen et al., 2020; Schroff et al., 2015; Zbontar et al., 2021; Bardes et al., 2022), contrastive PDE methods aim to leverage equation coefficients (Lorsung and Farimani, 2024), physical invariances (Zhang et al., 2023b), or Lie point symmetries (Mialon et al., 2023; Brandstetter et al., 2022b) to define differences in PDE dynamics that can be organized in a latent space. Another approach in PDE pretraining follows observed in-context learning and emergent behavior in LLMs (Wei et al., 2022; Brown et al., 2020; Radford et al., ) to design neural PDE solvers that are capable of following prompted PDE examples to forecast unseen dynamics (Yang et al., 2023a; Chen et al., 2024).
A more straightforward pretraining method focuses on directly training neural solvers to transfer to new PDE dynamics (Goswami et al., 2022; Chakraborty et al., 2022; Wang et al., 2022). This approach has also been scaled by training neural solvers with large and diverse training sets to characterize its transfer behavior (Subramanian et al., 2023), as well as shown to be generally more effective over other pretraining strategies (Zhou et al., 2024). As a step toward large-scale modeling, more principled training approaches have been proposed to learn PDE dynamics across diverse physics at scale. Recent work has proposed a combinatorial neural operator that learns different dynamics as separate modules (Tripura and Chakraborty, 2023), embedding separate PDEs to a common space to do multi-physics prediction (McCabe et al., 2023), incorporating denoising with scalable transformer architectures while training across diverse PDE datasets (Hao et al., 2024), and using a unified PDE embedding to align LLMs across PDE families (Shen et al., 2024).
@@
Masked reconstruction is a popular technique popularized by the language processing (Devlin et al., 2018) and vision (Dosovitskiy et al., 2020; Xie et al., 2021; He et al., 2021) domains to pretrain models for downstream tasks. Masked modeling is a broad field that spans many masking strategies, architectures, and applications (Li et al., 2024b); this ubiquity is attributed to scalability and architecture breakthroughs (Vaswani et al., 2023) that allow meaningful context to be learned through masked reconstruction (Cao et al., 2022). In the field of neural PDE solvers, masked pretraining has initially been explored as a method to pretrain neural solvers directly (Chen et al., 2024). However, we take a separate approach by investigating if models can understand physics through masked self-supervised learning and how these latent representations are manifested. After validating this learning method, we seek to understand if this knowledge can be used to benefit common PDE tasks.
Situating our Contribution
To frame our contribution within these past works, we draw comparisons with broader deep learning efforts to pretrain large encoders through self-supervision, such as BERT (Devlin et al., 2018) or CLIP (Radford et al., 2021; Ramesh et al., 2022), to be used in downstream tasks. For many reasons (e.g. data availability, architectures, pretraining strategies, compute resources), an equivalent work does not currently exist for PDEs. However, we hope to advance this research thrust to train general-purpose physics encoders, which could be of benefit to the PDE community. Specifically, pretrained encoders are often leveraged to accelerate model development and training by outsourcing certain architecture modules to pretrained models. Furthermore, having standardized, encoded representations of diverse PDEs can help in interfacing with conventional ML pipelines, whether it be allowing LLMs to interface with PDE data or diffusion backbones to condition on PDE inputs.
Figure 1: We investigate learning diverse PDE dynamics with masked autoencoders (MAE) and using learned representations to benefit various downstream tasks. (Masked Pretraining) An encoder is trained on unmasked patches of spatiotemporal PDE data, while a decoder reconstructs true data from latent encodings and learned mask tokens. (Supervised Fine-tuning) Pretrained encoders can be used to quickly regress equation coefficients or predict key PDE features. (Conditional Time-stepping) Neural solvers can achieve higher accuracy predictions through conditioning on MAE encodings. (Conditional Super-resolution) SR models can also benefit from conditioning on MAE encodings, using a discretization inversion ($D^{-1}$) and neural operator (Yang et al., 2023b) to predict high-resolution physics.
@@ 3 Methods
We describe our main methods in Figure 1. We first pretrain an encoder and decoder to reconstruct masked inputs. This pretraining objective can have a few benefits. Masking destroys inherent biases present in the data; especially at high masking ratios, this forces models to learn very general latent representations. Lastly, masked modeling does not assume any prior physics or leverage features specific to a particular PDE, making the approach applicable to a wide variety of physics.
After pretraining, we evaluate using the learned representations for three downstream tasks after discarding the decoder. The first task is to regress or classify PDE features, such as predicting coefficient values or boundary conditions of a PDE sample. This can be done by directly appending a linear head to the pretrained encoder. The second task is to improve the performance of neural solvers by conditioning on an encoded representation of input physics. Both the encoded and original representations of the data are passed to the neural solver. Lastly, we consider improving super-resolution performance by passing an up-sampled input to a neural solver and conditioning on an encoded representation of the low-resolution PDE input.
3.1 Masked Pretraining for PDEs
@@
We adapt the popular Masked Autoencoder (MAE) approach (He et al., 2021; Xie et al., 2021; Feichtenhofer et al., ) to train ViT models (Dosovitskiy et al., 2020; Arnab et al., 2021) to reconstruct 1D and 2D PDE data. Data are partitioned into non-overlapping patches before a random subset of these patches is sampled to be masked. The masked patches are omitted from the encoder input; the encoder embeds only the unmasked patches through a series of transformer blocks, which allows for larger encoders and faster training at large masking ratios (He et al., 2021). The encoded patches are then recombined with mask tokens according to their position in the PDE solution. Positional embeddings are added again to preserve positional information before the latent encodings and mask tokens are decoded. The decoder can be shallower and narrower than the encoder because it is discarded in downstream tasks (He et al., 2021), which further reduces training costs. The decoded tokens are projected into the PDE space through a linear layer before reshaping the patches to the input dimension. Lastly, the output is compared to ground truth PDE data through a MSE loss.
Within this general framework, the primary considerations affecting performance and compute are model size, masking ratio, and patch size. We study the effects of these parameters on the resulting model’s reconstruction performance. In general, we find that increasing model size requires more data to be effective. Furthermore, increased masking ratios decrease the reconstruction performance but may lead to better learned representations. Lastly, decreasing patch size improves reconstruction performance at the cost of increased compute. For additional details and hyperparameters we direct readers to Appendix A.
3.2 Lie Point Symmetry Data Augmentation
Table 1: Error at different augmentation probabilities
Aug. Prob. Error
0.00 2.48e-03
0.25 1.24e-03
0.50 1.17e-03
0.75 1.27e-03
1.00 1.37e-03
@@ 3.2 Lie Point Symmetry Data Augmentation
To emulate a larger pretraining dataset and aid in generalization, we investigate using Lie augmentations during pretraining, an approach that follows the use of data augmentations in vision or video pretraining (He et al., 2021; Xie et al., 2021; Feichtenhofer et al., ). Given a PDE, one can derive its Lie symmetries as a set of transformations $\{g_{1},\dots,g_{i}\}$, each with a variable $\epsilon_{i}$ that modulates the magnitude of the transformation. At training time, we apply $g_{i}$ sequentially, each with a randomly sampled $\epsilon_{i}$ to randomly augment PDE samples. This augmented PDE sample could represent a solution that has been shifted in space, time, or magnitude, among other transformations, but still propagates dynamics according to the original PDE. For a more detailed discussion of Lie point symmetries for PDEs, we refer the reader to Olver (1986) and Mialon et al. (2023). To understand the effects of Lie data augmentation on reconstruction accuracy, we apply randomly augment training samples with different probabilities, with 0 representing no augmentation and 1 representing always augmenting samples. The resulting validation MSE reconstruction losses are given in Table 1 after training on 1D KdV-Burgers data. We find that over-augmenting samples may shift the training distribution too much; furthermore, we generally find that the best results use smaller augmentation magnitudes $\epsilon$.
@@
PDE data are often discretized on a mesh, which can be both irregular and of different resolutions. This can present a challenge for vanilla transformers, which use the same positional encodings regardless of variations in input sequence length. This is especially pronounced in 2D; flattening two spatial dimensions results in positional information needing to wrap around inputs, and as a result, the same positional embedding can represent very different points in space when given different discretizations. Indeed, many approaches have been proposed to adapt ViT models to inputs of varying resolutions or patch sizes (Beyer et al., 2023; Tian et al., 2023; Dehghani et al., 2023; Fan et al., 2024). However, to avoid changes to the ViT backbone, we consider different strategies to accommodate varying input lengths. In particular, we consider the setup where inputs of shape ($n_{t},n_{x}$) or ($n_{t},n_{x},n_{y}$) can vary along their spatial dimension ($n_{x},n_{y}$) by multiples of the patch size ($p_{x},p_{y}$). Snapshots in time, or time windows, are sampled with length $n_{t}$, which discretizes the PDE trajectories to a common time dimension.
Table 2: Error w/ different resolution strategies
Strategy Error
None 4.40e-03
Pad 3.81e-03
Interp. 3.23e-03
Token 1.89e-03
The simplest strategy would be to pad the inputs to the maximum sequence length (Pad). However, this does not address the fact that positional information becomes overloaded and only provides information about input length. To address this, we consider interpolating positional embeddings from the maximum sequence length to variable sequence lengths (Interp.). This would allow tokens to receive approximate positions; however, this would not be accurate for irregular grids. Lastly, to adapt to arbitrary grids, we consider using an embedder network, which could be a 1D or 2D CNN (depending on the spatial dimension), to project the spatial coordinate grid to a token (Token). We compare the validation MSE reconstruction loss of using different strategies in Table 2 after training on multiresolution KdV-Burgers data, observing consistent improvement by using more sophisticated methods for embedding spatial discretizations. In 1D, multiresolution pretraining is significantly easier to learn, so we use the Interp. strategy for simplicity. In 2D, the use of tokenized spatial grid information becomes more important, so we use the Token strategy for 2D multiresolution experiments.
4 Experimental Setup
4.1 PDEs and Datasets
We describe a variety of PDEs used for masked pretraining and downstream evaluation. In 1D, we pretrain MAE models on the KdV-Burgers equation only, while in 2D we pretrain on the Heat, Advection, and Burgers equations simultaneously. In all PDEs, coefficients and forcing terms are randomly sampled to produce diverse dynamics within a dataset.